arthham

JEE Advanced 2016 · Probability (Class 12)

A computer producing factory has only two plants T1T_1 and T2T_2. Plant T1T_1 produces 20%20\% and plant T2T_2 produces 80%80\% of the total computers produced. 7%7\% of computers produced in the factory turn out to be defective. It is known that P(computer turns out to be defective given that it is produced in plant T1)=10P(computer turns out to be defective given that it is produced in plant T2)P(\text{computer turns out to be defective given that it is produced in plant } T_1) = 10P(\text{computer turns out to be defective given that it is produced in plant } T_2), where P(E)P(E) denotes the probability of an event EE. A computer produced in the factory is randomly selected and it does not turn out to be defective. Then the probability that it is produced in plant T2T_2 is

Answer: C. 7893\frac{78}{93}

Watch this solved on the board The full walkthrough, written out a line at a time with narration explaining each step as it appears. Pause anywhere and ask why a line is there. Free to watch

Or read the worked solution below. Jump to the steps ↓

Written solution

Given: P(T1)=20100,P(T2)=80100\text{Given: } P(T_1) = \dfrac{20}{100}, \quad P(T_2) = \dfrac{80}{100}

P(D)=7100,P(D∣T1)=10 P(D∣T2)P(D) = \dfrac{7}{100}, \quad P(D\mid T_1) = 10\,P(D\mid T_2)

To find: P(T2∣Dc)\text{To find: } P(T_2 \mid D^{c})

P(T2∣Dc)=P(T2) P(Dc∣T2)P(Dc)P(T_2 \mid D^{c}) = \dfrac{P(T_2)\,P(D^{c} \mid T_2)}{P(D^{c})} Let P(D∣T2)=x⇒P(D∣T1)=10x\text{Let } P(D\mid T_2) = x \Rightarrow P(D\mid T_1) = 10x P(D)=P(T1)P(D∣T1)+P(T2)P(D∣T2)P(D) = P(T_1)P(D\mid T_1) + P(T_2)P(D\mid T_2) 0.2(10x)+0.8x=0.070.2(10x) + 0.8x = 0.07 2.8x=0.07⇒x=0.072.8=0.025=1402.8x = 0.07 \Rightarrow x = \dfrac{0.07}{2.8} = 0.025 = \dfrac{1}{40} P(D∣T2)=140,P(D∣T1)=10x=14P(D\mid T_2) = \dfrac{1}{40}, \quad P(D\mid T_1) = 10x = \dfrac{1}{4} P(D)=0.2×14+0.8×140P(D) = 0.2 \times \dfrac{1}{4} + 0.8 \times \dfrac{1}{40} =0.05+0.02=0.07 ✓= 0.05 + 0.02 = 0.07\ \checkmark P(Dc)=1−P(D)=1−0.07=0.93P(D^{c}) = 1 - P(D) = 1 - 0.07 = 0.93 P(T2∩Dc)=P(T2) P(Dc∣T2)=0.8×0.975=0.78P(T_2 \cap D^{c}) = P(T_2)\,P(D^{c}\mid T_2) = 0.8 \times 0.975 = 0.78 P(T2∣Dc)=0.780.93=7893=2631P(T_2 \mid D^{c}) = \dfrac{0.78}{0.93} = \dfrac{78}{93} = \dfrac{26}{31} Answer: 7893=2631⇒option (3)\text{Answer: } \dfrac{78}{93} = \dfrac{26}{31} \Rightarrow \text{option (3)}

More from Probability (Class 12)

All 38 Probability (Class 12) questions →